The total cost of six books,five pencils and seven sharpeners is rs 115 and that eight books ,ten pencils and fourteen sharpeners is Rs 190,then which of the following article`s cost can be found uniquely?(a)book (b)pencil(c)sharpeners(d)none of these ...please give me answer clearly

Grade:10

1 Answers

manna
7 years ago
lets say b=books,p=pencils,s=sharpeners
according to question
6b+5p+7s=115-------------------------1
8b+10p+14s=190-------------------2
 
in eq 2 divide both side by 2
4b+5p+7s=95------------------3
f
substract both eq 1 and 3
 
we will get 2b=20
b=10
mean single book cost is 10 rs.
answer is book